OpenStudy (anonymous):

where do i graoh the fraction 2/5? A. between 2 and 3 B. between -1 and 0 C. between -3 and -2 S. between 0 and 1

OpenStudy (anonymous):

2/5 = 4/10 = 0.4

OpenStudy (anonymous):

so d?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks!